# TinyGraphEstimator: Adapting Lightweight Language Models for Graph Structure Inference

### v2

@@ 1 INTRODUCTION
To address this gap, we investigate whether small transformer-based language models can infer structural graph parameters directly from serialized graph representations. The models operate on deterministic graph format - edge list - and are tasked with predicting graph-theoretic measures from these structured inputs. This setup allows us to examine whether compact language models can generalize reasoning patterns over relational structures without relying on large-scale model capacity or specialized graph algorithms.
We evaluate several small, resource-efficient models on this task and further adapt them using the Low-Rank Adaptation (LoRA) fine-tuning method. The results show that while the models exhibit limited capability in the zero-shot setting, fine-tuning leads to consistent and substantial improvements across all structural measures. These findings demonstrate that compact language models, when properly adapted, can acquire reasoning skills for graph-structured data while maintaining high computational efficiency.
@@ 2 RELATED WORK
Recent advances in large language models (LLMs) have demonstrated impressive performance in tasks requiring symbolic reasoning, numerical computation, and pattern generalization [Brown et al., 2020, Bubeck et al., 2023]. Studies have shown that transformer-based architectures can internalize relational structures [Wang et al., 2025] and perform graph-related reasoning when provided with suitable prompts or structured inputs. Nevertheless, most prior work focuses on large-scale models with hundreds of billions of parameters, whose reasoning capabilities emerge from scale rather than explicit adaptation. In contrast, we investigate whether small language models - under 4B parameters - can infer precise graph-theoretic quantities. This shifts the focus from emergent reasoning in massive models to lightweight, interpretable reasoning in compact architectures.
A growing body of work explores how language models can represent or reason about graphs through textual or hybrid encodings. GraphGPT [Tang et al., 2024] and related studies encode adjacency information as token sequences, enabling LLMs to reconstruct or classify graph structures. Other works have investigated relational reasoning benchmarks [Hu et al., 2025, Agrawal et al., 2024], showing that LLMs can approximate certain structural statistics with appropriate context. Our approach builds on this foundation but emphasizes the systematic evaluation of structural parameter inference - quantitative graph descriptors such as density, transitivity, and global efficiency - using small models specifically adapted for this purpose.
@@
In contrast to prior work that primarily investigates either graph representation learning or large-scale emergent reasoning, our research unifies both directions by systematically testing the graph reasoning capacity of small, adaptable language models. By combining structured graph representations with efficient fine-tuning, our approach provides new insights into how compact LLMs internalize relational structure and approximate graph-theoretic computations.
@@ 3 PROPOSED SOLUTION
The proposed approach evaluates and enhances the ability of small language models to infer a wide range of graph-theoretic parameters directly from structured graph representations. Each model is prompted to predict a structured set of key structural properties that collectively describe its topology and complexity, based on the graph’s edge list representation. The inferred parameters include the detailed degree-based statistics such as density, minimum degree, mean degree, maximum degree, and degree standard deviation. In addition to these fundamental attributes, we evaluate the models’ ability to estimate higher-order structural metrics including the total number of triangles, average clustering coefficient, transitivity, average shortest path length, diameter, chromatic number, and global efficiency. Collectively, these parameters capture both local and global aspects of graph organization, providing a comprehensive benchmark for assessing reasoning over graph structures.
@@ 3.2 Dataset Construction
To the best of our knowledge, no existing dataset directly supports the task of mapping graph structures to their quantitative parameters. Therefore, we constructed a balanced synthetic dataset of connected graphs generated from three canonical random graph models: Erdős–Rényi (ER), Barabási–Albert (BA), and Watts–Strogatz (WS). The ER model produces graphs where each possible edge occurs independently with a fixed probability, capturing purely random connectivity [Erdös and Rényi, 1959]. The BA model generates scale-free networks through preferential attachment, emphasizing hub formation and heavy-tailed degree distributions [Barabási and Albert, 1999]. The WS model interpolates between regular lattices and random graphs, combining high clustering with short path lengths [Watts and Strogatz, 1998].
@@
Graphs are stored accompanied by metadata that capture key structural properties. The dataset consists of 1,200 graphs for training and 120 graphs for testing, evenly distributed across the three models. This design provides a well-controlled yet structurally varied benchmark for assessing the generalization and reasoning abilities of lightweight language models in inferring global and local graph parameters.
@@ 3.4 Validation Protocol
We evaluate model outputs with a schema- and tolerance-aware validation pipeline that mirrors the training prompt format while remaining model-agnostic. For each sample, the script loads the graph data and gold metadata and normalizes it to a canonical textual form. A structured prompt then instructs the model to return only a JSON object matching a fixed schema of given fields (nodes, edges, density, degree statistics, triangles, clustering, transitivity, path lengths, diameter, chromatic number, global efficiency). Inference is run on a base model or a base model augmented with LoRA adapters (if provided), with JSON-constrained decoding via lm-format-enforcer [Gat, 2024] to prevent malformed outputs. The first JSON object in the response is extracted and strictly checked against the schema.
@@ 4 RESULTS
Across all models, accuracy remains stable across different graph types, from sparse to dense topologies, indicating robust generalization and resilience to variation in degree distribution and connectivity. In summary, fine-tuning enables small, LoRA-adapted language models to perform precise quantitative reasoning over graph-structured data, serving as reliable, high-precision estimators of graph-theoretic properties.

### v3

@@ 1 INTRODUCTION
To address this gap, we investigate whether small transformer-based language models can be adapted to estimate structural graph parameters from textual representations of graphs. In our setup, each graph is encoded in a deterministic edge-list format, which serves as a textual proxy for its topology. The models are trained to predict key graph-theoretic quantities from these structured inputs, enabling us to assess whether compact language models can approximate graph reasoning patterns when fine-tuned on explicit structural data, rather than relying on built-in algorithmic mechanisms or large model capacity.
To this end, we systematically evaluate several small, resource-efficient models on the task and fine-tune them using the Low-Rank Adaptation (LoRA) method. The results show that while these models exhibit limited capability in the zero-shot setting, fine-tuning leads to consistent and substantial improvements across all structural measures. These findings demonstrate that compact language models, when properly adapted, can acquire effective reasoning behavior over graph-structured data while maintaining high computational efficiency.
@@ 2 RELATED WORK
Recent advances in large language models (LLMs) have demonstrated impressive performance in tasks requiring symbolic reasoning, numerical computation, and pattern generalization [Brown et al., 2020, Bubeck et al., 2023]. Studies have shown that transformer-based architectures can internalize relational structures [Wang et al., 2025, Schmitt et al., 2021, Ying et al., 2021] and perform graph-related reasoning when provided with suitable prompts or structured input [Huang et al., 2024, Guo et al., 2023, Guo et al., 2025]. Nevertheless, most prior work focuses on large-scale models with hundreds of billions of parameters, whose reasoning capabilities emerge from scale rather than explicit adaptation. In contrast, we investigate whether small language models - under 4B parameters - can infer precise graph-theoretic quantities. This shifts the focus from emergent reasoning in massive models to lightweight, interpretable reasoning in compact architectures.
A growing body of work explores how language models can represent or reason about graphs through textual or hybrid encodings. These approaches encode adjacency information as token sequences, enabling LLMs to reconstruct or classify graph structures [Tang et al., 2024, Chen et al., 2025, Podstawski, 2025]. Other works have investigated relational reasoning benchmarks [Hu et al., 2025, Agrawal et al., 2024], showing that LLMs can approximate certain structural statistics with appropriate context. Our approach builds on this foundation but emphasizes the systematic evaluation of structural parameter inference - quantitative graph descriptors such as density, transitivity, and global efficiency - using small models specifically adapted for this purpose.
@@
In contrast to prior work that primarily investigates either graph representation learning or large-scale emergent reasoning, our research unifies both directions by systematically testing the graph reasoning capacity of small, adaptable language models. By combining structured graph representations with efficient fine-tuning, our approach provides new insights into how compact LMs internalize relational structure and approximate graph-theoretic computations.
@@ 3 PROPOSED SOLUTION
The proposed approach evaluates and enhances the ability of small language models to infer a wide range of graph-theoretic parameters directly from structured graph representations. Each model operates on undirected, unweighted graphs whose topology is defined entirely by edge connectivity, and is prompted to predict a structured set of key structural properties that collectively describe graph topology and complexity. The inferred parameters include detailed degree-based statistics such as minimum degree, mean degree, maximum degree, and degree standard deviation, along with the global density metric describing overall connectivity. In addition to these fundamental attributes, we evaluate the models’ ability to estimate higher-order structural metrics including the total number of triangles, average clustering coefficient, transitivity, average shortest path length, diameter, chromatic number, and global efficiency. Taken together, these parameters capture both local and global aspects of graph organization, providing a comprehensive benchmark for assessing reasoning over graph structures.
@@ 3.1 Models Selection
For reference, we also evaluate two state-of-the-art LLMs - GPT-4.1 [OpenAI, 2024, Achiam et al., 2024] and DeepSeek-V3.2-Exp (Non-Thinking Mode) [DeepSeek, 2025, Liu et al., 2025] - to benchmark our compact models against the current frontier of LLM capability. While these systems exhibit strong general reasoning and broad-domain adaptability, our fine-tuned small models achieve superior accuracy on graph-structural inference, highlighting the efficiency and specialization benefits of lightweight adaptation.
@@ 3.2 Dataset Construction
To the best of our knowledge, no existing dataset directly supports the task of mapping raw graph structures to such a wide range of quantitative parameters. Therefore, we constructed a balanced synthetic dataset of connected graphs generated from three canonical random graph models: Erdős–Rényi (ER), Barabási–Albert (BA), and Watts–Strogatz (WS). The ER model produces graphs where each possible edge occurs independently with a fixed probability, capturing purely random connectivity [Erdös and Rényi, 1959]. The BA model generates scale-free networks through preferential attachment, emphasizing hub formation and heavy-tailed degree distributions [Barabási and Albert, 1999]. The WS model interpolates between regular lattices and random graphs, combining high clustering with short path lengths [Watts and Strogatz, 1998].
@@
Graphs are stored accompanied by metadata that capture key structural properties. Each graph is represented as an edge list, where every line specifies a pair of connected node identifiers. The dataset consists of 1,200 graphs for training and 120 graphs for testing, evenly distributed across the three models. This design provides a well-controlled yet structurally varied benchmark for assessing the generalization and reasoning abilities of lightweight language models in inferring global and local graph parameters.
@@ 3.4 Validation Protocol
We evaluate model outputs with a schema- and tolerance-aware validation pipeline that mirrors the training prompt format while remaining model-agnostic. For each sample, the script loads the graph data and gold metadata and normalizes it to a canonical textual form. A structured prompt then instructs the model to return only a JSON object matching a fixed schema of given fields (density, degree statistics, triangles, clustering, transitivity, path lengths, diameter, chromatic number, global efficiency). Inference is run on a base model or a base model augmented with LoRA adapters (if provided), with JSON-constrained decoding via lm-format-enforcer [Gat, 2024] to prevent malformed outputs. The first JSON object in the response is extracted and strictly checked against the schema.
@@ 4 RESULTS
Figure 1: Comparison of Large and Fine-tuned Small Language Models on Graph-Structural Inference. The figure compares average performance across R2 Accuracy and NRMSE Accuracy. Compact, fine-tuned small models substantially outperform state-of-the-art large models on graph-structural reasoning tasks, demonstrating the effectiveness of efficient, domain-adapted fine-tuning.
@@
In addition to the fine-tuned small models, we evaluated two state-of-the-art proprietary systems to contextualize the relative performance of our compact architectures. As illustrated in Figure 1, the small models substantially outperform these large-scale counterparts across both evaluation metrics. While GPT-4.1 and DeepSeek-V3.2-Exp achieve mean R2 Accuracies of 55.1% and 56.0%, and Mean NRMSE Accuracies of 28.5% and 39.4%, respectively, the fine-tuned small models reach over 94% and 78% on average for the same metrics. This consistent performance margin underscores the effectiveness of targeted, parameter-efficient adaptation over sheer model scale, demonstrating that compact models can achieve high reasoning precision when optimized for structural inference.
The fine-tuned small models maintain stable accuracy across diverse graph types - from sparse to dense topologies - demonstrating strong generalization and resilience to variations in degree distribution and connectivity. Overall, LoRA-based adaptation enables small language models to reason precisely and interpretably over graph-structured data, establishing them as efficient, high-precision estimators of graph-theoretic properties.
@@ 5 DISCUSSION
Beyond the presented benchmark, the proposed approach has practical implications for the development of lightweight reasoning agents and embedded AI systems. By enabling small language models to infer and interpret graph-structural properties, TinyGraphEstimator can support tasks such as knowledge graph analysis, network diagnostics, and symbolic planning under resource constraints. These capabilities highlight the potential of compact, fine-tuned models as efficient reasoning modules within larger autonomous or decision-support architectures.